rxcklesskaisher:

The graph below shows the price, y, in dollars, of different amounts of granola, x, in pounds: A graph titled Granola Prices shows Number of Pounds on x axis and Price in dollars on y axis. The scale on the x axis shows numbers from 0 to 12 at increments of 2, and the scale on the y axis shows numbers from 0 to 48 at increments of 8. A straight line joins the ordered pairs 0, 0 and 12, 48 Which equation best represents the relationship between x and y? y = x + 4 y = 8x y = 4x y = x + 8

rxcklesskaisher:

@jhonyy9

rxcklesskaisher:

@QuestionCoveBot

jhonyy9:

A straight line joins the ordered pairs 0, 0 and 12, 48 (0,0) and (12,48) when x= 0 result y=0 when x = 12 result y = 48 how you get from x=12 the y= 48 ?

rxcklesskaisher:

12x 4

jhonyy9:

perfect

rxcklesskaisher:

yes

jhonyy9:

questions ?

rxcklesskaisher:

yes

rxcklesskaisher:

what letter is it y=4x??

jhonyy9:

a,b,c,d, how you think from these ?

rxcklesskaisher:

C

jhonyy9:

y = x + 4 y = 8x y = 4x y = x + 8

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher C \(\color{#0cbb34}{\text{End of Quote}}\) yes ofc

rxcklesskaisher:

ok thanks

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher ok thanks \(\color{#0cbb34}{\text{End of Quote}}\) anytime yw.